_ _ _x_ _ =_ _ _ _ using the numbers 1 through 9 only once make this equation true
Answer by Edwin McCravy(20054)

There are 7 different solutions:

1.   138×42 = 5796
2.   157×28 = 4396
3.   159×48 = 7632
4.   186×39 = 7254
5.   198×27 = 5346
6.   297×18 = 5346
7.   483×12 = 5796

Edwin